OpenStudy (anonymous):

What is the solution to the system of equations? 2x+y=-3 x-2y=-4 ( , )

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

x = 2y -4

OpenStudy (anonymous):

All right so there are two methods. Substitution and combining. In this case, substitution would be easier

OpenStudy (anonymous):

its not the answer

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

you have to plug in x into the y equation

OpenStudy (anonymous):

which is y = 2x +3

OpenStudy (anonymous):

so, y = 2(2y-4) + 3

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

and for the x equation you plug in what you got for y

OpenStudy (anonymous):

and you get your two points

OpenStudy (anonymous):

i got 3, 4 ???

OpenStudy (anonymous):

@perl

OpenStudy (perl):

let me check with wolfram

OpenStudy (anonymous):

for x... x = 2(-2x-3)-4 x = -4x -10 5x = -10 x = -10/5 x = -2

OpenStudy (anonymous):

just do the same for the other one and you should get 1

OpenStudy (anonymous):

2(2) + y = -3 4 + y = -3 y = 1